Shift report 20030207 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Feb 07 2003 - 07:58:11 EST

  • Next message: Hironori Ito: "E-Log bug fix"
    ------------------------------------------------------------
    Date: Fri Feb  7 07:58:11 2003, Shift: 00:00-08:00
    Supervisor: kh, with: 
    Log book#14 pages 4 - 4
    
    Shift Summary:
    ------------------------------------------------------------
    No usable beam during the entire shift.
    At about 2:20, a store was achieved and cogged and steered to us with acceptable background.  MCR called and said that we were in good shape and that they would steer to STAR and then keep the store.  In the course of steering to STAR, our backgrounds went way up and then they lost the beam soon after.  Every attempted store after that died for some reason or the other.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Feb 07 2003 - 08:03:18 EST